Calculate the twisting couple of a hollow cylinder. 


1
Expert's answer
2021-03-08T08:30:23-0500

For a hollow cylinder of the same length land ofinner radius r1 and outer radius r2, Twisting couple of the cylinder


"C=\int_{r_1}^{r_2}\frac{2\pi N\theta }{l}x^3dx\\
C=\frac{\pi N\theta}{2l}(r_2^4-r_1^4)"